OpenStudy (anonymous):

What is the volume of the prism? A triangular prism is shown. The triangular base has a base of nine yards and a height of six yards. The height of the prism is sixteen yards. two hundred eighty-eight cubic yards four hundred thirty-two cubic yards five hundred seventy-six cubic yards eight hundred sixty-four cubic yards

OpenStudy (anonymous):

@austinL

OpenStudy (mathmate):

Hint: Area of a triangle = \( \frac{1}{2} base * height \) and area of a prism = area of base * height

OpenStudy (austinl):

 I believe this is what it is.^^ \[V = (\frac{1}{2}bh)\times l\]